# Evidence of Memory Effects in the Dynamics of Two-Level System Defect Ensembles Using Broadband, Cryogenic Transient Dielectric Spectroscopy

### v1

@@ Appendix C Dielectric Response under Different Drive Lengths
Figure 9: Electric field amplitude calibration at the sample plane. a, HFSS simulation of the electric field distribution in the waveguide at 3 GHz and 5 GHz. The simulation applies an input power of 1 W at the SMA port to excite the dominant mode. The snapshot is taken for both frequencies when the oscillating electric field reaches its maximum amplitude at the sample plane. b, Frequency-dependent electric field strength averaged over the sample plane and time (blue solid line). A scaling factor is applied to the pulse (red solid line) to achieve a uniform drive amplitude across the band, reducing the amplitude in regions with higher intrinsic field strength. The resulting average field profile is flattened (blue dashed line). c,d, Comparison of transient spectra for a photoresist-coated sapphire sample before (c) and after (d) field calibration. The calibrated spectrum shows reduced amplitude in previously over-driven regions and yields a more uniform interrogation of the TLS ensemble across the full spectral range. A short horizontal white line is added to indicate pulse bandwidth $\Delta$f.
Appendix D Electric Field Amplitude Calibration
@@
The waveguide design ensures that the transmitted signal is relatively constant over a bandwidth of approximately 2.8-6 GHz (Fig. 2b). However, the samples under test are coupled in the middle of the waveguide, which can be frequency-dependent as illustrated in Fig. 9. We leverage HFSS simulations to predict the electric field amplitude at the sample plane over our frequency bandwidth (blue dot and lines in Fig. 9b). To compensate for this variation, we apply a frequency-dependent scaling factor (red solid line in Fig. 9b). The impact of this calibration is evident when comparing the transient spectra of a photoresist-coated sapphire sample before (Fig. 9c) and after applying the correction (Fig. 9d). Significantly, most of the key features in the data remain. Still, such calibrations can ensure uniform electric field densities as a function of frequency and may prove critical in future experiments to deduce properties of individual TLS defect ensembles.
Appendix E Measurement Setup
@@
The pulses generated from the RFSoC passes through a room-temperature amplifier (Minicircuits CMA-83LN+) and a series of attenuators at various temperature stages of a Bluefors LD400 dilution refrigerator. The signal reaches the waveguide input port at base temperature (typically $<$ 10 mK) and interacts with the sample (Fig. 10). The output signal from the waveguide is collected using a readout SMA pin that then passes through a cryogenic circulator, HEMT (High Electron Mobility Transistor) amplifier, and two room-temperature amplifiers (Minicircuits CMA-83LN+), before returning to the RFSoC. A detailed wiring diagram is provided in Fig. 10.
@@ Acknowledgements.
Bert Harrop graciously diced the devices at Princeton. Simon Agnew helped perform the aluminum oxide deposition using an atomic layer deposition tool in William Scheideler’s lab at Dartmouth. We also acknowledge stimulating conversations with Kater Murch, Miles Blencowe, Geoffroy Hautier, Chandrasekhar Ramanathan, and Valla Fatemi. Startup funds from the Thayer School of Engineering at Dartmouth primarily supported this work. While the scientific focus of this work is quite distinct from existing funding, M.F. would like to express gratitude for the support of science within his group through the DARPA Young Faculty Award No.  D23AP00192 and NSF PHY-2412555, and hopes that projects like this continue to be made possible by grants from the US federal government. The views and conclusions in this manuscript are those of the authors. They should not be interpreted as representing the official policies, expressed or implied, of DARPA, the NSF, or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein.

### v2

@@ Appendix D Dielectric Response in Transmission and Reflection
Figure 10: Electric field amplitude calibration at the sample plane. a, HFSS simulation of the electric field distribution in the waveguide at 3 GHz and 5 GHz. The simulation applies an input power of 1 W at the SMA port to excite the dominant mode. The snapshot is taken for both frequencies when the oscillating electric field reaches its maximum amplitude at the sample plane. b, Frequency-dependent electric field strength averaged over the sample plane and time (blue solid line). A scaling factor is applied to the pulse (red solid line) to achieve a uniform drive amplitude across the band, reducing the amplitude in regions with higher intrinsic field strength. The resulting average field profile is flattened (blue dashed line). c,d, Comparison of transient spectra for a sapphire sample with AlOx deposit (same sample as Fig. 3e with a change in mounting configuration). We show the spectrum before (c) and after (d) field calibration. The calibrated spectrum shows reduced amplitude in previously over-driven regions and yields a more uniform interrogation of the ensembles of TLS defects across the full spectral range. A short horizontal white line is added to indicate pulse bandwidth $\Delta$f.
Appendix E Electric Field Amplitude Calibration
@@
The waveguide design ensures that the transmitted signal is relatively constant over a bandwidth of approximately 3-6 GHz (Fig. 2b). However, the samples under test are coupled in the middle of the waveguide, which can be frequency-dependent as illustrated in Fig. 10. We leverage HFSS simulations to predict the electric field amplitude at the sample plane over our frequency bandwidth (blue dot and lines in Fig. 10b). To compensate for this variation, we apply a frequency-dependent scaling factor (red solid line in Fig. 10b). The impact of this calibration is evident when comparing the transient spectra of a sapphire sample with AlOx deposit before (Fig. 10c) and after (Fig. 10d) applying the correction. Significantly, most of the key features in the data remain. Still, such calibrations can ensure uniform electric field densities as a function of frequency and may prove critical in future experiments to deduce properties of individual ensembles of TLS defects.
Appendix F Measurement Setup
@@
The pulses generated from the RFSoC pass through a room-temperature amplifier (Minicircuits CMA-83LN+) and a series of attenuators at various temperature stages of a Bluefors LD400 dilution refrigerator. The signal reaches the waveguide input port at base temperature (typically $<$ 10 mK) and interacts with the sample (Fig. 11). The output signal from the waveguide is collected using a readout SMA pin that then passes through a cryogenic circulator, HEMT (High Electron Mobility Transistor) amplifier, and two room-temperature amplifiers (Minicircuits CMA-83LN+), before returning to the RFSoC. A detailed wiring diagram is provided in Fig. 11.
Appendix G Qualitative Comparison to Numerical Simulations
The purpose of the numerics is to qualitatively reproduce the key experimental observation that the Fourier spectra exhibit broad sidebands for short pulses, which progressively sharpen as shown in Fig. 4 in the main text. To this end, we model the system using a driven-spin model with a finite-duration pulse, which provides a simplified framework for interpreting the interference patterns observed in the TLS ring-down and understanding how the applied drive shapes the dynamics. During this interval, the system is expected to exhibit Floquet-like physics, with the dynamics governed by the quasienergy spectrum and associated sidebands. After the end of the pulse, signatures of the prior Floquet dynamics may persist in the post-pulse evolution, manifesting as coherent oscillations or ringdown features. To support this interpretation, we briefly summarize the Floquet formalism and illustrate the emergence of sidebands using numerical simulations. We first briefly summarize the formalism following [72, 73]. Consider a time-periodic Hamiltonian $H(t)$ with period $T$,
$H(t+T)=H(t),$ (20)
and define the Floquet frequency $\Omega=\frac{2\pi}{T}$.
Floquet theorem states that solutions to the time-dependent Schrödinger equation are as follows.
$i\frac{d}{dt}|\psi(t)\rangle=H(t)|\psi(t)\rangle$ (21)
can be expressed as
$|\psi_{\alpha}(t)\rangle=e^{-iE_{\alpha}t}|u_{\alpha}(t)\rangle,$ (22)
where the Floquet mode $|u_{\alpha}(t)\rangle$ is periodic,
$|u_{\alpha}(t+T)\rangle=|u_{\alpha}(t)\rangle,$ (23)
and $E_{\alpha}$ is the quasi-energy of Floquet.
Expand $|u_{\alpha}(t)\rangle$ into Fourier harmonics,
$|u_{\alpha}(t)\rangle=\sum_{m=-\infty}^{\infty}e^{-im\Omega t}|u_{\alpha}^{m}\rangle,$ (24)
leading to the eigenvalue equation in the extended space
$\sum_{m}\mathcal{H}_{nm}|u_{\alpha}^{m}\rangle=E_{\alpha}|u_{\alpha}^{n}\rangle,$ (25)
where
[…]
For a monochromatic drive
$H(t)=H_{0}+Ve^{i\Omega t}+V^{\dagger}e^{-i\Omega t},$ (27)
the Floquet Hamiltonian $\mathcal{H}$ is block-tridiagonal:
$\mathcal{H}=\begin{pmatrix}\ddots&\vdots&\vdots&\vdots&\iddots\\ \cdots&H_{0}-\Omega&V&0&\cdots\\ \cdots&V^{\dagger}&H_{0}&V&\cdots\\ \cdots&0&V^{\dagger}&H_{0}+\Omega&\cdots\\ \iddots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}.$ (28)
Figure 12: Numerical simulation of collective excitation dynamics and Floquet quasi-energies in a driven two-spin system ($N=2$). a, Time-domain evolution of the collective population $\langle\sigma^{+}\sigma^{-}\rangle$ obtained by solving the Lindblad master equation for a square cosine pulse with dipole–dipole coupling $J=50\leavevmode\nobreak\ \text{MHz}$, drive amplitude $A=100\leavevmode\nobreak\ \text{MHz}$, pulse duration $\tau=100\leavevmode\nobreak\ \text{ns}$ (horizontal dashed line), collective dissipation rate $\Gamma=2.0\leavevmode\nobreak\ \text{MHz}$, and bare transition frequencies $\epsilon_{1}=3.5\leavevmode\nobreak\ \text{GHz}$ and $\epsilon_{2}=4.5\leavevmode\nobreak\ \text{GHz}$ (vertical dashed lines). b, Amplitude of the corresponding Fourier spectra of the time-domain response for the entire evolution as a function of drive frequency, revealing pronounced V-shaped spectral features that converge towards the bare TLS frequencies. c, Floquet quasi-energies (first four) computed during the driving period. Pronounced features appear near the driving frequencies that lead to long-lived ring-downs observed after the pulse is off in a.
Figure 13: Numerical simulation of a driven, interacting four-spin system ($N=4$) using the Lindblad master equation. The plots show a single disorder realization of an all-to-all connected spin network with randomly assigned bare frequencies and dipole-dipole couplings, drawn from uniform distributions $\epsilon\in[3.0,5.0]$ GHz and $J\in[-50.0,50.0]$ MHz, respectively. The system is subject to a collective radiative decay rate $\Gamma=1.0$ MHz and a drive amplitude $A=400$ MHz. a, Time evolution of the collective TLS population $\langle\sigma^{+}\sigma^{-}\rangle$ as a function of drive frequency for pulse durations $\tau=20$ (i), 50 (ii), and 200 (iii) ns. Tree-like ring-down structures emerge, becoming increasingly sharp with longer pulses. b, Corresponding Fourier spectra reveal broad sidebands for short pulses that progressively sharpen and converge toward the system’s bare eigenfrequencies as $\tau$ increases, giving rise to a distinct V-shaped pattern. The bare TLS frequencies are indicated by vertical dashed lines in a and b, while the pulse duration is marked by a horizontal white dashed line in a.
The dipole operator $\hat{P}(t)$ defines the polarization. The dipole response function is given by the retarded commutator.
$\chi(t,t^{\prime})=-i\theta(t-t^{\prime})\langle[\hat{P}(t),\hat{P}(t^{\prime})]\rangle,$ (29)
[…]
$\chi(t,t^{\prime})=\sum_{n,m}e^{-in\Omega t}e^{im\Omega t^{\prime}}\chi_{nm}(t-t^{\prime}),$ (30)
where the Fourier components satisfy the Dyson equation
$\chi_{nm}(\omega)=\chi_{nm}^{0}(\omega)+\sum_{k}\chi_{nk}^{0}(\omega)V_{km}\chi_{km}(\omega).$ (31)
Explicitly, the Fourier transform of the dipole response function is expressed via the Floquet Green’s functions, $G^{R}$, as
$\chi_{nm}(\omega)=\sum_{\alpha,\beta}\sum_{l}\frac{d_{\alpha\beta}^{l}d_{\beta\alpha}^{m-n-l}}{\omega-(E_{\alpha}-E_{\beta}+l\Omega)+i\eta},$ (32)
[…]
$\omega=E_{\alpha}-E_{\beta}+l\Omega$ (33)
determine the resonance frequencies.
When the system is driven by a periodic square pulse of duration $\tau$, then turned off at $t=\tau$, the system freely evolves with no drive for $t>\tau$. The post-pulse time evolution contains oscillatory terms at frequencies set by quasi-energy differences $E_{\alpha}-E_{\beta}$ and sideband shifts $l\Omega$, weighted by the residues of these poles in $\chi$. These produce ring-down oscillations, whose decay rates reflect the imaginary parts of $E_{\alpha}$.
For a driven two-level system with detuning $\epsilon$ and external field amplitude $A$, the time-dependent Hamiltonian is given by
$H(t)=\frac{\epsilon}{2}\sigma_{z}+A\cos(\Omega t)\sigma_{x},$ (34)
Here, the periodic drive induces dressed states with quasi-energies
$E_{\pm,m}\approx\pm\frac{\epsilon}{2}+m\Omega,$ (35)
which hybridize near multi-photon resonances $\epsilon\approx n\Omega$.
The effective coupling strength of $n$ photons is proportional to $2AJ_{n}\left(\frac{2A}{\Omega}\right)$ where $J_{n}$ is a Bessel function of the first kind. This effective coupling changes the size of the avoided crossing near multi-photon resonance. These avoided crossings form characteristic V-shaped structures near resonances in the $(\epsilon,\Omega)$ plane, where the arms of the V correspond to bare energy levels shifted by multiples of $\Omega$, and the vertex marks the opening of a gap due to multi-photon coupling. The tree-like structures observed in the ringdowns arise from the interference of multiple sideband emissions. Unlike the case of a single driven two-level system, our system consists of an ensemble of interacting TLSs coupled via dipole-dipole interactions. As a result, each eigenstate of the bare, time-independent Hamiltonian becomes dressed by the external drive, leading to more complex interference effects that manifest as beatings in the observed signal.
Although the discussion above focuses on a single two-level system for illustrative purposes, our numerical simulations consider systems with $N=2$ and $N=4$ coupled spins, incorporating weak dipole-dipole interactions, which are three orders of magnitude smaller than the individual TLS frequencies. Using this model, we observe collapse and revival patterns that qualitatively resemble the experimental observations. The Fourier transform of the simulated dynamics reveals characteristic V-shaped features in the spectral domain, converging at the bare eigenfrequencies of the coupled system. Dissipation is modeled through a collective radiative decay channel using the Lindblad master equation, which may be justified in our setup due to the presence of a broadband waveguide mediating the system–environment interaction. Hence, including interactions and collective dissipation, the Hamiltonian becomes [19]
$\hat{H}(t)=\sum_{j=1}^{N}\hat{H}_{0,j}-\hat{\mathbf{P}}\cdot\mathbf{E}(t)+\sum_{i<j}J_{ij}\hat{\sigma}_{x}^{(i)}\hat{\sigma}_{x}^{(j)}.$ (36)
The dynamics is governed by the Lindblad master equation.
$\frac{d\rho}{dt}=-i\left[\hat{H}(t),\rho\right]+\Gamma\left(2\hat{S}_{-}\rho\hat{S}_{+}-\hat{S}_{+}\hat{S}_{-}\rho-\rho\hat{S}_{+}\hat{S}_{-}\right),$ (37)
where $\hat{S}_{\pm}=\sum_{j=1}^{N}\hat{\sigma}_{\pm}^{(j)}$ are collective lowering and raising operators. The assumption of collective decay is motivated by the broadband nature of the waveguide, which couples similarly to all TLSs.
Figure 12 shows numerical simulations of the collective excitation dynamics and Floquet quasi-energies in a two-spin system. Panel a displays the time evolution of the collective population, obtained by solving the Lindblad master equation with fixed dipole-dipole coupling, drive amplitude, pulse duration (marked by a horizontal dashed line), collective dissipation, and two distinct TLS frequencies (indicated by vertical dashed lines). During the drive, clear Rabi-Chevron patterns emerge, while following the pulse, pronounced interference and long-lived ring-down oscillations are observed. Figure 12b shows the Fourier amplitude of the full time-domain simulation as a function of drive frequency. The V-shaped features converging toward the bare TLS frequencies reflect detunings between the drive and the individual transitions. These structures highlight the coherent excitation conditions that give rise to the long-lived ring-downs observed in panel a. Figure 12c displays the first four Floquet quasi-energies computed during the driving period. The quasi-energy spectrum exhibits characteristic dips near the driving frequencies, which correspond to resonances that enhance coherent excitation trapping. These resonances manifest as long-lived ring-down dynamics in the time domain, indicating that the system temporarily stores energy in quasi-stationary Floquet states with suppressed decay, thus prolonging the coherent oscillations observed after the pulse ends.
Figure 13 presents numerical simulations of a driven, interacting four-spin network under a finite-duration pulse. In the experiment, the emitted signal is measured via homodyne detection of the in-phase (I) and quadrature (Q) components, which are proportional to the collective dipole radiation from the system. This output field corresponds to the collective lowering operator $\sigma^{-}_{\text{tot}}=\sum_{i}\sigma^{-}_{i}$, so the measured power reflects the expectation value $\langle\sigma^{+}_{\text{tot}}\sigma^{-}_{\text{tot}}\rangle$, directly connecting to the simulated observable $\langle\sigma^{+}\sigma^{-}\rangle$. The upper panels show the time-domain evolution of this collective excitation as a function of drive frequency for increasing pulse durations, where pronounced interference patterns emerge in the ring-down following the pulse, as shown in Fig. 13a. Figure 13b displays the corresponding Fourier spectra, which reveal how the spectral content of the ring-down evolves with increasing pulse duration. For short pulses, the abrupt temporal cutoff introduces large energy uncertainty, resulting in broad sidebands. As the pulse duration increases, the drive becomes more spectrally selective, allowing coherent oscillations among near-resonant modes to build up over time. This temporal coherence leads to narrower sidebands in the Fourier spectrum, which progressively sharpen and approach the bare transition frequencies of the system. This sharpening gives rise to the observed V-shaped structure in the frequency domain that we observe experimentally.
The theoretical framework used here shows good qualitative agreement with the experimental data, capturing key features such as prolonged ring-downs that sharpen with increasing pulse duration, leading to more defined spectral features. This is particularly notable given the general assumptions underlying the model, especially in our TLS framework, where driving is incorporated within the standard tunneling model alongside dipole–dipole interactions. These encouraging results motivate future work aimed at a more detailed theoretical and numerical investigation of the atomistic origin and universal dielectric response of the ensembles of TLS defects probed by our broadband transient spectroscopy technique.
@@ Acknowledgements.
Bert Harrop graciously diced the devices at Princeton. Simon Agnew helped perform the aluminum oxide deposition using an atomic layer deposition tool in William Scheideler’s lab at Dartmouth. Sofia Abrego and Chloe Buschmann helped review and polish the manuscript. We also acknowledge stimulating conversations with Kater Murch, Miles Blencowe, Geoffroy Hautier, Chandrasekhar Ramanathan, and Valla Fatemi. Startup funds from the Thayer School of Engineering at Dartmouth primarily supported this work. While the scientific focus of this work is quite distinct from existing funding, M.F. would like to express gratitude for the support of science within his group through the DARPA Young Faculty Award No.  D23AP00192 and NSF PHY-2412555, and hopes that projects like this continue to be made possible by grants from the US federal government. The views and conclusions in this manuscript are those of the authors. They should not be interpreted as representing the official policies, expressed or implied, of DARPA, the NSF, or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein.